Express the following into m/s and km/h
288km/h
2km/h Solution 6omin=1hr 3600s=1hr 1000m=1km : 24*1000/3600 =20/3=6.667m/s

Answer by Ezekiel(1)

3600s=1hr 1000m=1km 24*1000/3600=20/3=6.667m/s